Nickel Too Small

The correct size of a nickel is 21.21 millimeters. Based on that, the data can be summarized into the following table:


Too Small

Too Large

Total

Low Income

20

20

40

High Income

24

11

35

Total

44

31

75

c) If 5 children are chosen at random, the probability that at least one would draw the nickel too small is:

d) If 110 children are chosen at random, it would be unusual if more than [how many?] drew the nickel too small.
Answer #1

a)

proportion of all children that drew the nickel too small is: =44/75 =0.587

b)

P(X=5)=(nCx)px(1−p)(n-x)    =0.2556

c)

P(X>=1)=1-P(X<=0)=1-∑x=0x-1   (nCx)px(q)(n-x) =0.9996

d)

more than 81 drew the nickel too small (please try 82 if this comes wrong)


answered by: gavin